OpenStudy (anonymous):

f(x) and g(x) if f (g(x))=18x-1

OpenStudy (anonymous):

f(x)=3x+2, g(x)=6x-13 f(x)=3x-5, g(x)=6x+2 f(x)=3x-2, g(x)=6x+13 f(x)=3x+5, g(x)=6x-2 These are the options. I am so lost.

OpenStudy (anonymous):

f(x)=3x+5, g(x)=6x-2

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

f(x)=3x+5,g(x)=6x-2

OpenStudy (anonymous):

Thanks @Isabel62

OpenStudy (anonymous):

your welcome mylove im glad to help